# A Fundamental Trade-off in Aligned Language Models and its Relation to Sampling Adaptors

### v1

@@ D.1 Examples of Sampling Adaptors
We recover ancestral sampling when ${f}(\cdot)={\mathbb{I}}(\cdot)$ and ${{C}\left({p}(\cdot\mid{{\boldsymbol{y}}_{<t}})\right)}={\overline{{{\Sigma}}}}$.
@@
We recover temperature sampling when ${f}({p}({{y}_{t}}\mid{{\boldsymbol{y}}_{<t}}))\propto{p}({{y}_{t}}\mid{{\boldsymbol{y}}_{<t}})^{\frac{1}{{\tau}}}$ and ${{C}\left({p}(\cdot\mid{{\boldsymbol{y}}_{<t}})\right)}={\overline{{{\Sigma}}}}$.
@@
We recover top-${k}$ sampling (Fan et al., 2018) when ${f}(\cdot)={\mathbb{I}}(\cdot)$ and
@@
We recover locally typical sampling (Meister et al., 2023b) when when ${f}(\cdot)={\mathbb{I}}(\cdot)$ and
@@
We recover $\eta$-sampling (Hewitt et al., 2022) when ${f}(\cdot)={\mathbb{I}}(\cdot)$ and
@@
Appendix E The trade-off in DPO-aligned models
@@ Results.
Appendix F Toy Experiment Distributions

### v2

@@ E.1 Examples of Sampling Adaptors
We recover ancestral sampling when ${\kappa}(\cdot)={\mathbb{I}}(\cdot)$ and ${{C}\left({p}(\cdot\mid{{\boldsymbol{y}}_{<t}})\right)}={\overline{{{\Sigma}}}}$.
@@
We recover temperature sampling when ${\kappa}({p}({{y}_{t}}\mid{{\boldsymbol{y}}_{<t}}))\propto{p}({{y}_{t}}\mid{{\boldsymbol{y}}_{<t}})^{\frac{1}{{\tau}}}$ and ${{C}\left({p}(\cdot\mid{{\boldsymbol{y}}_{<t}})\right)}={\overline{{{\Sigma}}}}$.
@@
We recover top-${k}$ sampling (Fan et al., 2018) when ${\kappa}(\cdot)={\mathbb{I}}(\cdot)$ and
@@
We recover locally typical sampling (Meister et al., 2023b) when when ${\kappa}(\cdot)={\mathbb{I}}(\cdot)$ and
@@
We recover $\eta$-sampling (Hewitt et al., 2022) when ${\kappa}(\cdot)={\mathbb{I}}(\cdot)$ and
@@
Appendix F The trade-off in DPO-aligned models
@@ Results.
Appendix G Toy Experiment Distributions